OpenStudy (anonymous):

(a) A space vehicle is launched vertically upward from the Earth's surface with an initial speed of 8.40 km/s, which is less than the escape speed of 11.2 km/s. What maximum height does it attain?

OpenStudy (aaronq):

\(\large \sf v_f^2=v_i^2+2a\Delta y\) \(\large \sf (0)^2=(8400~m/s)^2+2(-9.8~m/s^2)\Delta y\) \(\large \sf \Delta y=\dfrac{(8400~m/s)^2}{2(9.8~m/s^2)}=3600000~m=3600~km\)